OpenStudy (anonymous):

the question says: write the expression as the sine, cosine, or tangent of an angle. what the problem gives me is sin20degrees cosine10degrees+ cosine20degrees sine10degrees. pls help

OpenStudy (anonymous):

i will help

OpenStudy (jhannybean):

hint: \(\sin(x+y)=\sin(a)\cos(b)+\cos(a)\sin(b)\)

OpenStudy (jhannybean):

Now you figure out what your \(\mathbf x\) and \(\mathbf y\) are